### v1

# Information Seeking Using AI Assistants

Abstract
A good portion of a software practitioners’ day involves seeking and using information to support task completion. Although the information needs of software practitioners have been studied extensively, the impact of AI-assisted tools on their needs and information-seeking behaviors remains largely unexplored. To addresses this gap, we conducted a mixed-method study to understand AI-assisted information seeking behavior of practitioners and its impact on their perceived productivity and skill development. We found that developers are increasingly using AI tools to support their information seeking, citing increased efficiency as a key benefit. Our findings also amplify caveats that come with effectively using AI tools for information seeking, especially for learning and skill development, such as the importance of foundational developer knowledge that can guide and inform the information provided by AI tools. Our efforts have implications for the effective integration of AI tools into developer workflows as information retrieval and learning aids.
Index Terms:
AI-assisted tools, information seeking, software development, productivity
I Introduction
Artificial Intelligence (AI) has revolutionized how we think about software engineering. According to the most recent StackOverflow Developer Survey, 76.7% of developers are already using or plan to incorporate AI-assisted software development assistants into their workflows [1]. Developers use AI-assisted software tools in a variety of contexts and to complete a diversity of tasks, most often hoping for increases in their productivity [2].
A significant contributor to, or deterrent from, developer productivity is the time they spend seeking information to support the completion of their tasks [3, 4]. Over the years, we have acquired an in-depth understanding of the information needs developers have [5, 6, 7], the ways in which they attempt to meet their information needs [8], and the challenges they encounter in the process [9]. We also know that developers’ ability to meet information needs contributes to their ability to acquire and build expertise [10, 11].
Given the changing landscape in developer tooling, many studies have investigated the use and impact of AI-assisted tools on developer engagement and productivity [12, 13, 14, 15] . However, most of these studies have focused on AI-assisted tool use in the context of understanding, generating, or modifying source code, providing little insight into the role AI plays in developer information seeking and its impact. To address this gap, we conducted a mixed methods study to answer the following research questions:
RQ1
When, why, and how do developers use AI tools for information seeking?
RQ2
What impact does the use of AI tools for information seeking have on developer productivity?
RQ3
What impact does the use of AI tools for information seeking have on developer skill development?
To answer our research questions, we first administered a survey to better understand to what extent developers are engaging with AI tools in this context. Building on those insights, we conducted a series of 17 interviews to further contextualize those experiences. Our findings reveal that while AI tools can offer increased efficiency when seeking information, there are caveats to realizing the benefits, such as avoiding over-reliance and building the necessary expertise to appropriately and effectively use AI tooling to meet information needs.
II Related Work
The goal of our research is better understand information seeking with the availability of AI-assisted tools. Most relevant to our efforts investigations into the information needs and seeking behaviors of developers and human-centric concerns regarding the use of AI tools in software development.
II-A Information Needs of Developers
Previous studies have examined the information needs of developers during the software development process. One of the first efforts centered on developer information needs was conducted by Ko et al. [5], where they observed developers in their work to better understand information needs of collocated software teams. Also aiming to gain a broad understanding of developer information needs is the work done by Fritz et al. [8], who conducted a series of 11 interviews to explore the use of information fragments to answer developers’ questions. On the flip side, LaToza et al. [9] surveyed 179 professional developers regarding hard-to-answer questions about code to better understand information needs developers encounter challenges trying to fulfill.
Some prior efforts are more focused, with an interest in understanding more specific information needs. Breu et al. [16] analyzed information needs in bug reports by examining questions asked in six hundred bug reports from the Mozilla and Eclipse projects. Liu et al. [6] investigated API-related developer information needs on StackOverflow by annotating StackOverflow questions to APIs. Phillips et al. [7] explored information needs for integration decisions in the release process of large-scale parallel development by interviewing seven release managers. With the emergence of AI-assisted tools, our builds on prior efforts by providing insights into the ways in which these tools have changed software practitioners’ information-seeking behaviors.
II-B AI Tool Use in Software Engineering
Researchers have also studied the impact of integrating AI tools into software engineering processes. Zheng et al. [17] conducted a literature review and identified 123 papers focusing on large language models (LLMs) in software engineering, highlighting their utility for various development tasks. Most other works in human-centered AI-assisted tool use focus on understanding and supporting developer interactions. Barke et al. [18] investigated how developers interact with AI programming assistants like GitHub Copilot. They classified interaction modes into two types: acceleration mode, where developers know the next steps and use Copilot to speed up their work, and exploration mode, where developers are unsure and use Copilot to discover possible solutions. Liang et al. [19] conducted a large-scale survey to investigate the usability of AI programming assistants, focusing on the successes and challenges developers face when using these tools. Johnson et al. [20] developed the PICSE framework, which identifies key factors influencing engineers’ trust and usage of traditional and AI-assisted software tools. Their framework was derived from interviews with software practitioners. Given the concern regarding the impact of AI-assisted tool use on productivity, prior work has also investigated how we can measure developer productivity when using code completion tools like GitHub Copilot1[12]. Our work builds on these prior efforts by providing insights into how these developer interactions with AI tools influence not just task completion, but also their information-seeking behaviors and its impact on their productivity and skill development.
III Methodology
The goal of our study is to better understand information seeking behaviors when using AI tools to complete software development tasks. Below we describe the study we conducted towards this goal.
III-A The Survey
We designed a 20-minute survey administered through Qualtrics to engage developers in our research.  [^1] This research is approved under IRBNet #: [Blind for review]. Our survey was divided into four sections. The first section asked a series of demographic and background questions, such as current job role and years of programming experience. The design of this section was heavily influenced by the annual StackOverflow Developer Survey [21]. The next section asked questions about the specific AI-assisted tools they use, the frequency and purpose of use, and rationale behind their usage. We also asked questions about their information seeking behaviors with and without AI assistance, including the advantages and disadvantages for each. To assess the impact of AI tools on development tasks, we also asked respondents to evaluate how these tools influenced their approach to tasks like debugging, testing, implementing, and planning. In the final section of the survey, we included questions regarding the role of AI tool in learning and integrating new technologies, such as the impact on learning curves and the potential pitfalls. We also gave respondents the opportunity at the end of the survey to express interest in a follow-up interview (Section III-B). Prior to administering our survey, we piloted with 3 respondents. This helped ensure clarity and identify any issues before deployment. We excluded this data from our final dataset. The survey instrument is available in our supplemental materials.
Our goal was to recruit developers with various levels of experience, technical skills, job roles, and team sizes. Therefore, we advertised our survey in both virtual and physical settings. We promoted our efforts through the personal LinkedIn and Twitter accounts of all authors, as well as through internal mailing lists. We also reached out to developer communities in the Washington D.C. Metropolitan area, where most of the authors are based. We also used snowball sampling by encouraging participants to share the survey with their networks. In total, we received 310 responses, with 173 participants expressing interest in participating in a follow-up interview.
Data Preparation
Given we distributed our survey through social platforms, there is a heightened risk of receiving invalid responses [22]. To mitigate the potential for analyzing invalid data, we filtered out responses that were completed in under three minutes or incomplete. This initial filtering left us with 168 responses. We further excluded responses that contained irrelevant content in their open-ended answers, as this suggests they may not have been giving due consideration to the survey. This resulted in our final dataset, containing 128 valid responses, which we used to report our findings.
TABLE I: Age Groups of Survey Respondents
Age Group No. Respondents
18-24 years old 44
25-34 years old 73
35-44 years old 6
45-54 years old 3
55-64 years old 2
65 years or older 1
TABLE II: Professional Programming Experience
Among Survey Respondents
Years of Experience No. Respondents
0-2 years 36
2-5 years 53
5-7 years 21
7-10 years 10
More than 10 years 8
Active UsersOccasional Users002020202040404040606060608080808010010010010083.683.683.683.616.416.416.416.4Percentage (%)
Figure 1: Usage of AI Tools Among Survey Respondents
TABLE III: Job Titles of Survey Respondents
Job Title No. Respondents
Developer (full stack, frontend or backend) 62
Research & Development Role 15
Data Scientist or Machine Learning Specialist 14
Data or Business Analyst 13
Project Manager 5
DevOps Specialist 4
Security Professional 4
Other 3
Designer 3
Senior Executive (C-suite, VP, etc.) 2
QA or Test Engineer 1
System Administrator 1
Cloud Infrastructure Engineer 1
TABLE IV: AI-Assisted Tools Used by Survey Respondents
Tool No. Respondents
ChatGPT 118
GitHub Copilot 62
Gemini 51
Microsoft Copilot 38
Tabnine 8
Claude 4
Other 8
Respondents
In our final dataset of responses (128), the age of our survey respondents ranged from 18 to over 65, median being 25-34 [I] with of 4 years of experience [II]. Prior research suggests that developers with less experience spend more time seeking information and learning as they transition from novices to experts [10] [11] [23], making this demographic particularly valuable for our study. Majority of our survey respondents actively use AI tools [1] such as ChatGPT and GitHub Copilot [IV] in their work. In addition to formal education, a majority of our survey respondents also used books and online resources such as forums and courses to learn programming. We interviewed a subset of these respondents, depending on their interest and availability and report on their demographics in  V.
Data Analysis
To analyze our survey, we first mapped each survey question to our research questions. We categorized experience levels of participants into five different groups, as outlined in Table II. For closed ended questions, we ran descriptive statistics and applied Fisher’s exact test with Bonferroni correction to identify any correlations. We also report on frequencies to supplement our findings. For open-ended responses, we used thematic summaries to report findings due to the small sample size of valid and useful responses. In the following section, we only report significant findings.
III-B The Interviews
To supplement our survey findings, we conducted interviews over Zoom to gather detailed insights into participants’ information-seeking practices, experiences with AI-assisted tools, and their impact on productivity and skill development. This helped us obtain a variety of perspectives and allowed for immediate follow-up questions for clarification. Each interview began with a discussion of participants’ background in software development, their experiences with AI-assisted tools, and how these tools fit into their workflow. Topics covered included the impact of AI tools on productivity, skill development, and problem-solving strategies. We also asked questions regarding their roles and responsibilities, the typical activities in their jobs, and their reliance on other developers or software tools. This includes what they do use, or avoid use of, AI tools for when information seeking and their satisfaction with the interactions. We elicited specific examples of their use to better understand aspects such as workflow integration and problem solving approaches. Lastly, we inquired about the impact team dynamics and productivity.
We piloted the interview protocol following best practices with one author and two participants to identify and correct any issues. The final interview protocol is available in our supplemental materials. We continued interviewing participants until we reached theoretical saturation, where no new themes or insights emerged from the interview [24]. Table V details the background of our interview participants.
TABLE V: Interview Participant Background
Participant Years of Experience Job Title
P1 2.5 Software Developer
P2 3 Software Developer
P3 2 Technical Analyst
P4 3.5 Data Analyst
P5 3 Software Developer
P6 2.5 Software Developer
P7 1 Business Analyst
P8 1.5 Software Developer
P9 3 Software Developer
P10 0.5* Software Developer
P11 4 Research Engineer
P12 1.5 Data Engineer
P13 2.5 Software Developer
P14 2 Software Developer
P15 5 Software Developer
P16 6 Software Developer
P17 18 Software Engineering Lead
* has background in instructional design and writing data analysis
software prior to current role
Analysis
To analyze our interview data, we used qualitative coding and thematic analysis, beginning with the creation of an initial codebook based on questions asked during the participant interviews. In the first iteration, all authors independently labeled an interview transcript using the initial codebook and then collectively discussed their findings, which led to the emergence of a new set of codes. After agreement on the codes and how they should be used, we incorporated the emergent codes to finalize the codebook. Following this, the first author applied the final codebook to label the remaining transcripts.
To ensure the rigor and validity of our efforts, we invited an external auditor to review a subset of raw coded segments from the interview transcripts along with the codebook. We asked the auditor to confirm that the inferences we made from the data were sound. We discussed and clarified any concerns and updated code descriptions when deemed necessary. After validating the codes, the first author reviewed all coded segments under each code and documented the overarching themes as they emerged. Finally, we mapped the codes to our research questions and iteratively identified any overlapping themes. Given our goal was to uncover themes present in the data, rather than their frequency,and following best practices for interview analysis [25], we do not report any Inter-rater Reliability (IRR).
IV Results
The goal of our efforts are to uncover when, why, and how developers are using AI tool to support their information seeking. Below we discuss our findings regarding the considerations developers may make when deciding to engage with AI tools to meet their information needs.
IV-A Information Seeking with AI tools (RQ1)
AI-Assisted Information Seeking
Our findings suggest that developers are often using AI tools to support their information seeking (Figure 2 and Figure 3), with the majority reporting using AI tools at least half of the time in their work for support (104) and seeking information (93). However, according to the experiences reported in our interviews, their goals when doing so are most often broad and aimed at understanding the necessary considerations for completing their task.
5 (Always)4321 (Never)002020202040404040414141413131313132323232171717171111# Participants
Figure 2: Frequency of AI Use for Support
5 (Always)4321 (Never)002020202040404040222222224040404025252525383838383333# Participants
Figure 3: Frequency of AI Use for Information Seeking
According to our survey respondents, the most notable benefits of using AI tools for information seeking include increased efficiency (92), more personalized responses (75), and programming language-agnostic support (43). One of our interview participants noted that using AI tools requires active engagement since “the AI is driving it” and this can go “on a tangent or in the right direction”. A few participants also discussed how using AI tools helps reduce the “burden” imposed on other developers when asking for help, as they then have to “spend time on our issues”.
Most participants reported relying on AI tools when trying to understand best practices, discover new libraries or solutions or explore trade-offs between different libraries and implementations. In this context, one participant stated “even if it can’t solve a complex coding problem, it can often point me to best practices for how to solve that problem.”[P14] Many participants noted that AI tools can help find links to relevant tutorials for in-depth learning while some participants reported that it can provide tailored explanations for new concepts using analogies or equivalents from their existing knowledge.
“And then I would also ask the GPTs or AI assistant for relevant links or relevant tutorials, so that I can actually learn about this process in general, and then apply that methodology or principle in my specific work.”[P8]
Participants in our study also cited AI tools as being useful for providing clear answers to generic conceptual questions, which can also aid in identifying keywords for further online searching and validation. Another common use case among participants when information seeking is to use AI tools to recall previous knowledge or explain code. Participants mentioned value in using AI tools for synthesizing relevant information from documentation, which includes providing boilerplate code, pinpointing highly specific issues or information. One participant shared a concrete example of this
”I can find pretty much any niche thing I need to know about AWS, whereas before, it required sifting through extensive documentation and numerous Amazon support pages.”[P15]
UX Issues
While participants in our study outlined numerous benefits to AI-assisted information seeking, concerns surfaced regarding the usability of AI tools. Most prominent were discussions regarding the ways in which the AI tool communicate the information. A major theme that emerged was around the conversational tone used by most AI-assisted tools. One participant discussed how this can undermine the utility of AI tools for tasks like code reviews.
“But for something like code review, it’s a bit riskier because error handling is not just a nice-to-have; it’s a must-have. While the AI correctly identifies many issues, it frames them as improvements rather than necessities. For example, it might suggest adding error handling or not hardcoding secrets, or it might recommend CSRF protection on endpoints as good improvements. However, these practices are actually mandatory.”[P15]
In addition, some participants cited that the tendency for AI tools to agree with the user also decreases their reliability as an information seeking tool. When recalling their own experiences, one participant stated, ”It provided the correct information initially but later changed its stance to align with what I was saying. So now I don’t know when to trust it.” Participants also reported issues with inappropriate information framing, finding the level of detail present in their responses to be ineffective, noting that “Sometimes they tend to give out too much information, sometimes too little.” We discuss more about the impact this has on information seeking when trying to learn new technologies in Section IV-C.
D*CA*R*CC*B*LNT*0020202020404040406060606080808080565656566666666632323232353535356868686881818181666666667575757554545454626262624747474745454545TasksCountTraditionalAI
Figure 4: Traditional vs. AI Preference
11footnotetext: Legend for task symbols: D = Debugging, CA = Code Analysis, R = Refactoring, $C$ = Code Comprehension, B = Brainstorming, LNT = Learning New Technologies.
Validation Techniques
While our findings point to consistent use of AI tools for information seeking, we found participants often only partially rely on information provided by AI tools, acknowledging the limitations of AI-generated solutions. Expressing sentiment shared by many participants, one participant stated, “Chat GPT 3.5 was providing a poor and incorrect implementation for network config files that were in YAML. ChatGPT3.5 also provided a partial answer that it scraped from Python docs regarding ArgParse library. I had to go to the actual docs to find it had a lot of missing context and prior steps in order to use the code it generated.” Prior work has emphasized concerns regarding the accuracy, privacy and security issues that can come with using AI tools  [20]. In addition to these concerns, our survey respondents indicated that when using AI tools to seek information they have concerns regarding decreases in self-directed learning and skill-building(76), privacy and security (66), and reduced information accuracy (60).
Given the skepticism around AI-generated information, our participants also discussed a variety of methods for validating the information provided. When it comes to information seeking endeavors that result in generated code or information used to modify their code, one common approach among our participants was unit testing. This involves running unit tests against any code modifications resulting from AI tool use to verify its functionality, check the output against expected results, and identify any issues or bugs.
Some participants noted using AI tools to support their validation efforts. In some cases, this meant using cross-validation, where they would compare responses from multiple AI assistants to identify conflicts or inconsistencies. For example, one participant stated that when one AI tool generates some code that needs an exhaustive set of test cases, “it’s especially helpful if you use multiple AI assistants for this” as it will help them “find a whole pile of issues”.
A common approach regardless of code being provided was iterative refinement, an AI-supported strategy where participants provide feedback to the AI assistant when the generated information is not satisfactory, modify prompts to obtain more accurate or appropriate information, and continuously refine the process until the desired outcome is achieved. They also rely on their own research and expertise, along with the expertise of others, to validate the information and confirm the accuracy and reliability of the AI-generated solutions. Or, as put by one participant, “there is always Google”. Another participant shared an experience they had with this kind of iterative solution refinement:
“I was trying to solve an error. And no matter how much I changed the point or how much I explained, it was still giving me the same solution. So I had to change my approach. I just went over to Google, and I entered whatever information I needed for that particular library to run. I went to StackOverflow, asked for different solutions and one of those solutions worked for me.”[P12]
Traditional Information Seeking
When discussing their use of AI tools, many participants described tasks, methods, and tools they rely on for information seeking when AI falls short. Many survey respondents indicated that AI tools are more useful than traditional methods for a number of tasks (Figure 4). However, interview participants often discussed the need for human-centric resources when unable to resolve the issue or complete the task with AI-generated information alone.
“So if I’m stuck, I go to other colleagues, but then, before going to my colleague or mentors for help, [I’ve been] using this [AI tool] as help to solve the problem.”[P13]
The most common resource mentioned for information seeking without AI assistance was StackOverflow, which participants cited using for context-specific tasks, such as determining the appropriate regular expressions for their use case or environment setup, in the hopes that other developers have encountered similar scenarios. Another common source participants mentioned using was Reddit, where they would seek general guidance, such as tool recommendations, from specific subreddits. We also found that when it comes to tasks that require specific information, knowledge, or context, or when working with specialized or niche libraries, our participants sought support directly from other developers or support teams on contract. When attempting to summarize the capabilities of AI tools for supporting information seeking in comparison to working with other developers, one participant stated, “It does it about as well as I would expect for a junior software engineer with a little bit more domain knowledge than I would expect from a a junior software engineer, but without a lot of the personality or experience, or trust that I would get from a junior software engineer.”
One participant highlighted the benefits of discussing their information needs with their colleagues, emphasizing that such interactions often lead to valuable, in-depth conversations and innovative ideas. This was especially the case when seeking information on coding practices and patterns for a large codebase, or working in niche spaces.
Policy & Practice
Regardless of the benefits, and perhaps due in part to the concerns, that come with using AI tools for seeking information, participants in our study highlighted several external factors that influence the impact and their use of AI tools in software development. One such factor is organizational restrictions that aim to reduce the risks associated with using AI tools, which may include limitations on using certain AI tools like Copilot or the public version of ChatGPT. Our findings suggest organizations may also have agreements with providers such as Microsoft , driven by concerns regarding protection of Personally Identifiable Information (PII). Legal protections offered by major AI providers like Microsoft and Google, which pledge to defend users in AI-related copyright lawsuits, may also influence the confidence and extent to which organizations adopt AI tools.
Participants also noted that their usage of AI tools for information seeking is heavily influenced by the adoption and promotion of AI within the organization. Some organizations are proactive in adopting and promoting AI use, aiming to utilize these tools responsibly and in a controlled manner, while others feel “it’s not worth your time”. However, this proactive stance can sometimes lead to an overshadowing of traditional engineering practices, resulting in less investment in non-AI related professional development.
“I’d say on the negative side, I feel like AI has taken over the industry so much that it’s pushed out some traditional engineering. And so it’s less motivating for employers to want to invest in learning a unique knowledge or unique tool… non technical people who run private companies [tend to] think that AI could just solve things. So there may not be as many professional development resources for anything non-AI related.”[P17]
IV-B Impact on Perceived Productivity (RQ2)
Most of our participants reported the use of AI tools for information seeking having a positive impact on their productivity. More specifically, participants cited time savings when seeking information for trivial, routine, or repetitive tasks. Most notably, AI tools support quicker finding and consolidation of relevant information and sometimes offer step-by-step guidance on how to make use of the information. Participants also mentioned AI tools reducing time spent blocked by obstacles, as AI tools offer quick solutions and guidance, allowing for efficient problem-solving and smoother project progression. For some participants, the time savings in these more trivial tasks also supports their ability to focus on more complex tasks.
“I use an AI assistive tool [when] staring at a blank screen and trying to figure out, okay, what is it that I need to do next? What component do I need to do? What do I need to start with right now?… One of the [AI] tools will break that down into smaller steps and help break the flow a bit to help me get back into the flow.”[P10]
In fact, despite the potential for new tooling to be disruptive to developer workflows, which can have a negative impact on productivity, participants from our survey and interviews indicated AI tools may not be as counter-productive when information seeking. One participant emphasized this distinction, stating, “The efficiency gained is worth the loss of flow. But for me, flow is less about thinking and more about implementing. The AI is doing the implementing, so the work I am doing is overcoming the roadblocks, not doing the easy work that flows.”
For most of our survey respondents, AI tools at the very least have no impact on their workflow (109), where respondents reported that their workflow was about the same when using AI tools as when they do not. For some respondents, mostly those with fewer years of development experience, using AI tools for information seeking actually improves their flow (78). Those who felt AI tools improved their flow, the rationale was most often not having to look at as many sources, being provided step-by-step guidance, and faster learning. As indicated by many participants, using AI tools supports getting “a lot of stuff in one place” rather than “searching multiple links” returned from, for example, from a Google search.
Measuring Productivity
While time savings was the most common metric of productivity among our participants, our efforts uncovered other ways developers think about productivity when using AI tools for information seeking. When considering time savings, this can be the time required to complete tasks or projects, debug and resolve issues, and make decisions. We also found that output quantity was another metric; this refers to the number of tickets or tasks completed per day or sprint, as well as publication output for research and development teams. One participant noted that they could “get a ticket that has been scoped for several days or a week done in a half an hour”. Another stated when they can close “[more] tickets [than usual] in a day” or “[multiple] feature development and a couple of bug fixes” then they know “the tool is assisting [me] in the right way.” Some participants also discussed other considerations that affect how they measure their productivity when using AI tools. For example, some participants emphasized the role of solution quality in their productivity, which includes things like frequency of issues arising and the number of code review iterations required.
We found that some productivity-related considerations had an impact on the decision to use an AI tool for information seeking to begin with. One such consideration is cost-effectiveness which is measured by the resources required (e.g., number of developers involved). As put by one participant, “ I would say, if we are able to finish this chat analysis project. Oh, my God! That would save thousands of person hours.”
Project and Team Impact
Our participants also highlighted the impact of using AI tools for information seeking beyond the individual developer. For some, there have been visible impacts to the projects they are working on when using AI tools. Participants cited AI assistants making it easier to parse large volumes of documentation to help pinpoint and resolve issues faster and providing code optimization strategies that speed up application. Some also discussed AI tools helping them clarify requirements and consider edge cases, leading to faster sprint completion times and more robust solutions. When discussing their own experiences, one participant described a concrete example of this kind of support:
“I needed to handle a bulk of requests from customers’ mobile devices. Initially, I assumed that the number of requests would not exceed a certain limit in a real-life scenario. However, there was an instance where this assumption was proven wrong, leading to a significant surge in requests. I quickly consulted with GPT, and we implemented a quick fix to resolve the issue that day. Since then, I’ve started considering potential edge cases more rigorously. I now regularly check with GPT, asking for potential edge cases based on my approach. The suggestions I receive help me anticipate potential issues, allowing me to proactively incorporate necessary adjustments into my design to avoid future problems.”[P6]
On the flip side, participants also discussed the negative impacts AI tool use can have. Most noticeable is the impact AI tools have on software teams. Our participants report reduced face-to-face interactions and organic knowledge sharing, which can impact team cohesion and collaboration. As described by one participant, some may be more inclined to use AI even when collaborators are an option:
“I would say one negative effect of it has been that we don’t have as much synergy with each other. And I think this also ties in sometimes to the fact that we’ve mainly gone to working remote. Because we’re no longer forced to bounce things off each other. I would say we’re less incentivized to bounce ideas off each other, ask for help on concepts, and I believe some of that organic conversation has gone missing. So for me … I happen to work with a very renowned scientist in that area. But, you know, in my mind, I’m more likely to go to like AI for help.”[P11]
Participants feel especially more comfortable using AI tools when they are less knowledgeable about something, as they may feel self-conscious about approaching someone more expert.
“We used to have different Slack channels for like Python help, or Excel, or web development. Those have kind of, I wouldn’t say gone silent, but they’re, if anything, mainly used for jokingly sharing a really bad answer from an AI. But at the end of the day I think it’s preferable for them to use AI instead of outing themselves not knowing a certain coding convention or making a really silly error, because it’s not anonymized on Slack.” [P11]
IV-C Impact on Skill Development (RQ3)
Using AI Tools to Learn
Many participants discussed how using AI tools when seeking information supports their learning and skill-building. Most often, participants mentioned the effectiveness of AI tools for filling in knowledge gaps and providing practice problems that support hands-on learning. For these participants, using AI tools in this way helps reinforce new concepts and improve problem-solving skills. In fact, the most frequently cited use for AI tools was explaining and clarifying concepts relevant to their tasks or goals. Participants often discussed an appreciation for AI tools’ ability to break down complex topics into more understandable parts. Our participants also reported frequently using AI tools for learning best practices, citing them as valuable resources for understanding industry standards and improving coding techniques.
AI-driven information seeking can also support the discovery and adoption of new solutions. An overwhelming majority of survey respondents reported using AI for learning new technologies (119). Participants in our study found AI tools useful for exploring innovative approaches and alternative methods for solving programming challenges, where they noted being able to get acquainted with new technologies without the overhead of seeking and learning about each individually. One participant noted being “more willing to use new tooling I don’t understand, barely understand, or rarely work with.” AI tools have also played a role in building other participants’ confidence in using new libraries or technologies, though some still reported discomfort in applying the knowledge gained from AI tools independently to practical scenarios. For one participant, this can be a product of overreliance, stating, “if you keep looking at ChatGPT for more and more alternate solutions, it just makes you lose your confidence.”
While many of our interviews suggest some usefulness of AI tools for learning new things, a high proportion of respondents from our survey indicated that they experience a higher learning curve when learning about and understanding new technologies using AI. In fact, for some participants, AI tools are not suited for supporting their learning, despite their ability to provide personalized responses. For some, the information presented by AI tools often provide inadequate levels of detail. As stated by one participant:
“The information is not typically presented in a pedagogically or intelligently designed way, like, you have to know what question to ask to be able to get an answer. But there is no pedagogy. There’s no there’s no like ”Oh, I can tell, based on this question that you’re just now learning this, and let me teach you in a way that I expect you to be able to retain this information or to give you more thorough context or links to resources that you can investigate on your own that are like good verifiable links” or anything like that. So yeah, I would not use it as a primary source of learning.”[P17]
Impact on Skill Development
When discussing the ways in which they feel AI tools impact their skill-building, many participants cited appreciating the ability to expand their knowledge and technical skillset more quickly with AI tools. However, there were also numerous concerns surrounding the skill building aspect of using AI tools for meeting information needs. From our survey, we found that most of our respondents have concerns regarding using AI tools for learning and skill-building because use in this context could lead to knowledge gaps (88), overreliance on AI tools (94), and poor problem solving skills (87).
One notable drawback that emerged from our interviews is the decrease in the ability to build creative problem-solving skills when engaging AI tools for meeting information needs. As indicated by participants, this can lead to more fundamental issues such as decreased ability to code from scratch and reduced understanding of the technologies being used.
“It’s like when we were doing calculations on our own. But then calculators came in. So we don’t calculate it in our mind anymore Whenever I come across a problem I just put it into GPT. I just don’t think about it on my own anymore, because I am relying a lot on AI tools… I just put it in there if I think it will help. It has greatly reduced my [ability to think independently].”[P7]
This includes diminished learning regarding use of traditionally fundamental software development tools and techniques such as command line, scripting, and debugging, where they can now “just throw the error, or whatever problem,…to ChatGPT and it will fix it.” This is exaggerated by the fact that the knowledge provided by AI tools is often limited to the specific question asked, often lacking the depth and breadth required for comprehensive understanding. In addition, iterative interactions are often required where the user must continuously refine their questions to receive useful answers. Some participants emphasized the challenge of not knowing what to ask or how to ask AI tools for information when learning about something new. As one participant noted, “it only covers the range of questions I have in my mind…it is like giving an answer to your question, but not delivering a session on one topic.”
While using AI tools can reduce creativity and critical thinking, some participants in our study also noted that AI-assisted information seeking can formulate a different kind of creative and critical thinking skillset. For example, according to some participants, correcting or refining an AI tool’s suggestions requires creativity, critical thinking, and problem solving skills. One participant made this notion concrete, stating:
“…and when I’m done I approach GPT for practice. So I’ll actually ask the GPT to give me [practice] problems that are, for example, I’ll take a component and I will ask GPT to give me [problems to solve] in different scenarios. So then I would … start thinking of [how to solve] them. And then I would ask it again to give me the solution when I’m done [figuring out] my own so that I can compare and think. It helps me think in a different way.”[P2]
This process can also foster creativity by sparking new ideas and approaches the developer may not have initially considered. A handful of participants reported that AI tools actually improve creativity, as they provide a foundation upon which developers can build and expand their ideas. As noted by one participant, “I think it promotes creativity. Because you may be introduced to new ways or new patterns of doing something…I think that it sometimes, by being too wordy, will like accidentally give people ideas that they otherwise wouldn’t have, or like expose people to ideas that they otherwise wouldn’t have. Yeah, I’d say it’s like a net positive win in the creativity department”[P17].
Our survey respondents also indicated that they use AI tools for both learning about (114) and integrating new technologies (113) in their work. However, from our interviews, we found there may be differences in the impact of using AI tools for learning about new technologies rather than integrating new technologies. One participant summarized this well, stating:
“I think AI tools can provide a temporary solution for overcoming a particular situation, but they can make it harder to truly understand the core knowledge or information that the technology or task involves. Since AI tools give us direct steps and solutions, we end up doing less research on our own, which can limit our knowledge and skills in the long run.”[P1]
Participants felt that while AI can expedite problem solving, it can also lead to a superficial understanding of new technologies if used for integration, which may be the reason why some participants reported feeling less confident and comfortable in using these technologies independently for learning. When discussing their use of AI tools for learning, one participant noted that AI “generally gives us an overview or brief about it, but not the complete information we need.”
V Discussion
Our findings indicate a shift in how developers are seeking information and building expertise in the age of AI assisted tooling. Below we discuss important insights into broader, existing concerns around the increasing integration of AI into software development.
V-A AI Tools Cannot (Always) Replace Developers
Two major, and related, concerns that have emerged with the rapid advancement and integration of AI tools in software development is that eventually nobody will have to learn to program anymore  [^2] https://www.techradar.com/pro/nvidia-ceo-predicts-the-death-of-coding-jensen-huang-says-ai-will-do-the-work-so-kids-dont-need-to-learn and AI will replace developers [26]. These concerns are likely heightened by the development of tools like ChatGPT and Gemini that provide human-like interactions. Furthermore, as our findings suggest developers may be engaging with each other less frequently when using AI tools, this raises important concerns regarding the future of collaborative software development; if developers are hyper-focused on AI tools facilitating their work, perhaps because they are too intimidated to talk to a teammate, this could greatly impact the ability for teams to coordinate heavier tasks, like integration, and to identify when teammates are duplicating efforts. While our findings suggest that there may be tasks that AI tools can perform more efficiently, and even effectively, than developers, they also emphasize how unlikely this future is.
Prior work has already emphasized the importance of social support in programming [27] and provided insights into the distinctions between AI and developers [28]. Our findings provide further evidence to debunk these myths by emphasizing the role of the developer in AI tools’ ability to support information seeking and meet their needs. For example, one theme throughout our findings was that an AI tool is only as effective as the interactions the developer has with it. This is the whole premise of efforts in prompt engineering that aim to understand how to get the best outcomes from AI tools [29, 30, 31]. But most importantly, our work amplifies how far AI tools are from being self-sufficient, relying heavily on developer knowledge and expertise for guidance towards relevant, high-quality information that can meet their goals and facilitate efficient task completion. This suggests that developers may be more often perceiving these tools as copilots that work in synergy with them rather than independently automating a handful of tasks, such as generating boilerplate code. Even then, according to our findings, there are some gaps that AI tools currently do not and can maybe never fill, such as trust, personality, and focused expertise, that make software development both fruitful and enjoyable.
V-B AI-Assisted Info Seeking Is Productive (Sometimes)
Another concern that has been raised in the software engineering community is regarding the negative impacts that AI tools can have on productivity and the ability to measure the impact it is having [12, 32, 33, 2]. Our findings suggest that developers perceive that using AI tools for information seeking can lead to significant and worthwhile productivity gains. In fact, despite any potential disruption to existing workflows when using AI tools, and the impact prior work suggests this may have [15, 34, 20], our findings suggest that the efficiency gains may be worth the loss of flow or even reduce the disruptions by reducing context switches that are often inherent in seeking information [35, 36]. In fact, we found that there are a range of scenarios where developer believe AI-assisted tools are increasing their productivity. Prior work has already emphasized the potential productivity gains from using AI-assisted tools for code completion [12]. Our findings suggest these potential productivity gains extend much beyond code completion into tasks such as consolidating large volumes of information, providing step-by-step guidance, filling in knowledge gaps, and dealing with the “blank page problem” of not knowing how to get started.  [^3] https://www.forbes.com/sites/forbesagencycouncil/2023/07/20/generative-ai-and-solving-the-blank-page-problem/
Our efforts also provide novel insights into the specific productivity-related considerations when measuring productivity and impact of AI tool use in information seeking. The most dominant theme across our findings was that of simplicity. As found in prior work focused on using AI tools for coding tasks, we found that information seeking with AI tools provides the biggest boost in productivity when the ask is simple. In the context of meeting information needs, this means avoiding questions that requires knowledge that cannot be easily accessed. For example, an AI tool can easily curate relevant content for a given tool, technology, or concept, which would save the time it would have taken for a developer to do it themselves. However, asking the AI to provide information to resolve a specific issue or complete a task can lead to wasted time, as well as less than favorable outcomes, given there may be context or insight not being considered by the AI.
V-C AI Tools Cannot (Alone) Support Skill Development
As suggested by prior efforts, and emphasized in our findings, AI tools are increasingly being adopted to increase developer efficiency when learning [37]. Our work provides detailed insights into the potential for AI tools to increase efficiency when learning and developing skills, which is how developers advance from novice to expert as programmers and in their workplace [10]. Research has investigated the ability to learn using AI tools [38, 39, 40]. However, most of this work has been situated in classroom settings where the methods and goals for learning are different than for developers, as our findings also emphasize that they most often acquire additional or new knowledge through the completion of their work tasks.
This important distinction points to a need to better understand and facilitate responsible use of AI tools in the context of building expertise. This is especially the case given the open-ended nature of most AI tools, where they can facilitate both task-oriented and goal-oriented information seeking. In our efforts, we found that AI tools are best suited for supporting the retrieval and consolidation of information, which can indirectly facilitate or potentially expedite learning or skill building. However, our findings also emphasize the importance of foundational knowledge in the effectiveness of AI tools in this context and the risks of over-reliance. This suggests a necessary shift in how and when AI tools should be used, promoting the use of AI tools when learning or understanding is the goal (e.g., what are the best practices for X?) within the context of a task (e.g., fix this bug) as opposed to using them to complete the task for them. In fact, our findings suggest if the goal becomes too close to the task (e.g., is this fix sufficient?), the nature of AI tools’ communication style can lead to an incorrect understanding of solution quality.
This also indicates a shift in the perception of important skills for developers. For example, many participants in our study felt it is becoming less important to exhaustively learn programming languages and more important to learn the fundamentals and best practices, as only when the necessary foundations are present can AI tools effectively and efficiently fill in the gaps. This also supports the ability to more effectively utilize AI tools as learning aids, as our findings indicate it may be difficult to use AI tools in this way without having the knowledge required for adequate prompting and iteration. These considerations are especially important given the impact our findings suggest appropriate (or inappropriate) use of AI tools can have on developer confidence in their knowledge and skills (and ability to apply them in practice).
Our findings also point to the potential for a future shift in where knowledge is found, stored, and accessed and the impact that may have on AI-assisted tools. Evidence of this potential shift can be found in prior work [41] as well as trends over the past couple of years on the StackOverflow Developer Survey, where the percentage of developers who frequently visit StackOverflow decreased from 2022 [42] to 2024 [43]. Our findings suggest this decrease may be happening due to an increase in use of AI tools for goals that developers would typically visit StackOverflow to fulfill. While on the surface this may not appear problematic, reduced interactions on these platforms will inevitably lead to fewer documented question and answer text which impact the ability for AI tools to learn from relevant, up-to-date information. This could also become an issue for StackOverflow and other knowledge repositories where the ability to scale becomes virtually non-existent.
VI Threats to Validity
Internal. Our study relies on self-reported data from surveys and interviews, which may be subject to memory bias. We asked participants to ground their response in past experiences to mitigate this threat.
External. Our sample was recruited through LinkedIn, X, and internal mailing lists, which might not be representative of the broader developer community. Developers who are not active on these platforms or in these networks might have different experiences with AI tools. To mitigate this, we also engaged with local developer communities and recruited participants through snowball sampling. Human-centered empirical studies are also prone to generalizability issues, due in part to concerns like sample size. However, the goal of our study is not to be generalizable, but rather transferable [44]. We ensure this by using a mixed-method approach where we supplement our survey findings with interviews. In our qualitative analysis, we followed best practices to ensure rigor. We did not report qualitative data using quantitative methods to prevent misinterpretation of our findings [45]. To further ensure the validity of our thematic analysis, we invited an external auditor to review our methodology and findings, providing an additional layer of validation.
Construct. As with any survey or interview, there is potential for misinterpretations questions by participants. To mitigate this, we piloted our survey and interview protocols with multiple participants to ensure clarity and refined questions as needed. We also provided definitions for key terms to ensure participants had a consistent understanding of the concepts being discussed.
VII Conclusion
In this paper, we describe our efforts exploring how information-seeking behavior has evolved in the era of AI tools. Based on data collected from a survey and set of interviews, we report on the kinds of information developers use AI tools to seek, challenges that come with using AI tools for information seeking, and the impact this has on developer productivity and skill development. Our work provides novel insights and implications regarding the importance of foundational knowledge in effective AI tool use, the potential for AI tools use to increase productivity, and best practices for AI tools as learning aids in software development.